# Think-at-Hard: Selective Latent Iterations to Improve Reasoning Language Models

### v1

@@ 6 Conclusion
This paper provides sufficient information to reproduce the reported results. All experiments were conducted using publicly available datasets together with open-source models and code. Appendix A details implementation aspects, including data selection, hyperparameters, and training procedures. To facilitate full reproducibility, we will release the code, configuration files, and model checkpoints upon publication.
@@ A.1.1 Training Recipe
As shown in Table 5, we follow the official setup in Open-R1 (openr1).We set maximum length to 8192 for Standard, TaH, and TaH+, and to 4096 for AlwaysThink due to its substantially higher memory usage during training.
@@
A.1.2 Latent Overthinking Analysis
We set up an oracle experiment to estimate the performance upper bound of our method. The oracle employs the DeepSeek-R1-Distill-Qwen-32B model as a dynamic label generator, replacing the MLP-based iteration decider. During the forward pass, we compare the token predictions from the label generator with those from the TaH backbone model. The model continues to iterate only when the top-1 predictions of these two models differ. Due to resource constraints and computational overhead, we evaluated the accuracy only on the first 100 samples from the MATH500 dataset.
A.1.3 Token Alternation Patterns
$\Pr\big(c_{i}^{(1)}>c_{\text{threshold}}\mid t_{i}=t\big),$
But 34.3% So 13.63%
Wait 12.17%
Therefore 8.95%
So 17.7% So 28.17%
Therefore 13.67%
But 4.89%
A.1.4 Routing Baseline Setup
For the routing baseline (Section 5), we route between MobileLLM‑R1‑360M and Qwen3‑1.7B, and between Qwen3‑0.6B and Qwen3‑4B. The routing endpoints are SFT‑trained under the same settings as the Standard baseline in Section 5, and the two routing pairs are configured to match average parameter sizes of 0.6B and 1.7B, respectively.
@@ A.2 Additional Experimental Results
A.2.1 Iteration Depth Beyond 2
@@
Hard Token Labeling. Previous works have proposed many methods to evaluate the hardness of each tokens in the training data, like through excess loss (lin2024rho; xie2023doremi), entropy (wang2025beyond; chen2023eeLLM) and prediction difference (fu2025r2r).
@@
Table 7: Performance comparison between TaH-2 and TaH-3 (maximum per-token iterations of 2 and 3, respectively). Iter.2 and Iter.3 denote the per-token percentages executing 2 and 3 iterations, respectively.
@@
A.2.2 Domain generalization
To assess cross-domain effectiveness, we train a 1.7B TaH model on the science subset of Open-R1. We exclude responses longer than 4,096 tokens, yielding 223M tokens in total; all other training settings follow Section 5. We then evaluate TaH against the Standard baseline on the GPQA-diamond benchmark (rein2023gpqa). TaH improves performance from 35.4 to 39.9, demonstrating the generalizability of our method.
A.2.3 Additional Latent Thinking Methods
Some latent thinking methods requires pre-training and uses base model other than Qwen3. We also compare with these methods, including Ponder (zeng2025pondering). Specifically, we adopt the released pretrained PonderingPythia-1.4B as the base model and perform SFT on the same training data. We observe that the fine-tuned model learns the stylistic patterns of the training data, but still underperforms substantially, which may be attributable to the limited capability of the PonderingPythia-1.4B backbone.
@@ A.2.3 Additional Latent Thinking Methods
Table 9: Per-token FLOPs and average iteration count for Standard, AlwaysThink, TaH and TaH+.
Standard AlwaysThink TaH TaH+
Param. Dataset FLOPs Iter. FLOPs Iter. FLOPs Iter. FLOPs Iter.
0.6B OlympiadBench 1.4e9 1.00 5.6e9 2.00 1.5e9 1.09 1.5e9 1.06
GSM8K 1.1e9 1.00 4.5e9 2.00 1.2e9 1.07 1.2e9 1.07
MATH500 1.3e9 1.00 5.3e9 2.00 1.4e9 1.05 1.4e9 1.06
AMC23 1.4e9 1.00 4.0e9 2.00 1.4e9 1.05 1.5e9 1.05
AIME25 1.5e9 1.00 5.8e9 2.00 1.6e9 1.05 1.6e9 1.06
A.2.4 Iteration Decider Analysis

### v2

@@ Acknowledgments
This paper provides sufficient information to reproduce the reported results. All experiments were conducted using publicly available datasets together with open-source models and code. Appendix A details implementation aspects, including data selection, hyperparameters, and training procedures. To facilitate full reproducibility, we release the code, configuration files, and model checkpoints at https://github.com/thu-nics/TaH.
@@ A.1.1 Training Recipe
We follow the official training setup of Open-R1 (Hugging Face, 2025) and use the Mixture-of-Thoughts dataset. For our main performance experiments (Section 6.2), we filter samples with output length exceeding 8K tokens from each category (math, code, and QA), then randomly sample 33K examples from each to form a balanced 100K training set. The filtered dataset contains 480M tokens, with 1% reserved for validation. For design choice exploration (Section 6.3), we use only the math subset filtered by 8K output length, resulting in 75K training samples. For 0.6B and 1.7B models, we use a maximum sequence length of 8192 tokens for all methods. For 4B models, we reduce the maximum sequence length to 4096 for Standard, TaH, TaH+, and 3072 for Ouro due to memory constraints. Detailed training hyperparameters are listed in Table 3.
@@
A.1.2 Baseline Setups
AlwaysThink. AlwaysThink uses the same architecture as TaH, except that it iterates twice at every token position and uses standard causal attention (attending only to the current iteration depth) instead of duo-causal attention.
Ouro. We implement Ouro architecture following its official design (Zhu et al., 2025), but fine-tune it on the Open-R1 dataset from Qwen-Base initialization, to align with all other methods. During fine-tuning, we adopt the entropy-regularized loss of Ouro. Since the original Ouro is trained from scratch on different data distributions, performance may differ from the original paper.
A.1.3 Parameter Breakdown
Table 4 reports the parameter breakdown of the Standard, TaH, and TaH+ methods. To offset the additional parameters introduced by TaH through LoRA and the iteration decider, we remove one layer from the LLM backbone, ensuring a fair comparison. In practical deployments, we recommend TaH+, which adds only about 3% additional parameters.
Table 4: Parameter breakdown of Standard, TaH, and TaH+. Counts are reported using M (million) and B (billion).
Param. Method Backbone LoRA Iter. Decider Total
0.6B Standard 596M 596M
TaH 580M 10M 5M 595M
TaH+ 596M 10M 5M 611M
1.7B Standard 1.72B 1.72B
TaH 1.67B 17M 18M 1.71B
TaH+ 1.72B 17M 18M 1.76B
4B Standard 4.02B 4.02B
TaH 3.92B 32M 70M 4.02B
TaH+ 4.02B 33M 70M 4.12B
A.1.4 Latent Overthinking Analysis Setup
We set up the oracle experiment in Table 14 to estimate the performance upper bound of our method. The oracle employs the DeepSeek-R1-Distill-Qwen-32B model as a dynamic label generator, replacing the MLP-based iteration decider. During each iteration, we compare the token predictions from the label generator with those from the TaH model. The model continues to iterate only when the top-1 predictions of these two models differ. Due to resource constraints and computational overhead, we evaluated the accuracy only on the first 100 samples from the MATH500 dataset, denoted as MATH100 throughout the paper.
@@ A.2 Additional Experimental Results
A.2.1 Performance on Math Benchmarks
Table 5: Accuracy comparison of different baselines across five benchmarks and two model sizes. Subscripts indicate improvement over Standard. The top two scores for each task and model size are highlighted in bold.
Method
Param. Benchmark Standard Routing SoftThink AlwaysThink TaH TaH+
0.6B AIME25 4.2 1.0 2.5 1.5 4.2 5.0
OlympiadBench 18.8 7.4 19.4 10.2 23.9 24.0
AMC23 23.4 10.9 24.1 15.6 32.5 30.6
MATH500 47.2 27.3 48.8 32.8 51.2 54.2
GSM8K 62.5 45.6 61.3 54.6 64.4 68.8
Average 31.2 18.5 31.2 22.9 35.2/+/+4.0 36.5/+/+5.3
1.7B AIME25 13.3 10.2 12.9 10.0 17.9 14.6
OlympiadBench 33.0 30.6 33.4 30.0 38.8 41.2
AMC23 42.2 42.2 43.1 42.5 48.4 51.2
MATH500 68.4 60.0 68.8 61.8 74.4 73.0
GSM8K 82.1 71.2 79.6 79.3 84.5 85.8
Average 47.8 36.8 47.6 44.7 52.8/+/+5.0 53.2/+/+5.4
4B AIME25 23.3 22.5 22.5 OOM 30.4 28.3
OlympiadBench 47.7 45.0 50.1 50.5 52.0
AMC23 62.8 60.9 64.1 70.3 70.6
MATH500 82.8 76.1 83.2 84.4 85.6
GSM8K 90.5 85.3 90.9 90.4 91.5
Average 61.4 58.0 62.2 65.2/+/+3.8 65.6/+/+4.2
Math Benchmark Evaluation. In this section, all models are trained on the math subset of Open-R1 and evaluated on math benchmarks (see Section A.1.1 for training details). We also add a Routing baseline, which selects a model from a candidate pair for each question. In our experiments, we use two pairs: (1) MobileLLM-R1-360M, Qwen3-1.7B, and (2) Qwen3-0.6B, Qwen3-4B. All candidate models are SFT-trained under the same settings. For each pair, the routing ratio is calibrated so that the average active parameter count matches our 0.6B and 1.7B targets, respectively. Table 5 reports results for 0.6B, 1.7B, and 4B backbones across five challenging math benchmarks. Even with strong Qwen3-Base initialization, existing approaches show limited effectiveness: AlwaysThink and routing methods fail to consistently outperform the standard baseline, while SoftThink yields only marginal gains. In contrast, TaH achieves stable improvements, with average gains of 4.0% (0.6B), 5.0% (1.7B), and 3.8% (4B) over Standard. TaH+, which adds less than 3% parameters, further improves to 5.3%, 5.4%, and 4.2%, respectively. For 0.6B and 1.7B, TaH achieves 8.1–11.3% gains over AlwaysThink, and TaH+ achieves 8.5–12.6% gains. AlwaysThink-4B is not evaluated due to out-of-memory during training.
Out-Of-Distribution (OOD) Performance. We further evaluated the zero-shot generalization capability of models trained solely on math datasets from the main paper. As shown in Table 6, TaH+ demonstrates consistent improvements not only on in-domain math benchmarks (MATH500, AMC23) but also on out-of-domain tasks like MMLU-STEM. This indicates that the thinking patterns learned by TaH+ on math problems are robust and transferrable to broader scientific reasoning tasks.
Table 6: Performance of math-only trained models (0.6B and 1.7B) on in-domain math benchmarks and the out-of-domain STEM benchmark (MMLU-STEM).
Param. Benchmark Standard SoftThink AlwaysThink TaH+
0.6 MATH500 47.2 48.8 32.8 54.2
AMC23 23.4 24.1 15.6 30.6
MMLU-STEM 51.6 51.4 42.6 56.3
Average 40.7 41.4 30.3 47.0
1.7 MATH500 68.4 68.8 61.8 73.0
AMC23 42.2 43.1 42.5 51.2
MMLU-STEM 70.8 70.6 63.8 73.7
Average 60.5 60.8 56.0 66.0
A.2.2 Real-world Efficiency
Setup. We investigate the real-world efficiency of different 1.7B models under our current implementation. All measurements were obtained on a single A800 GPU with a batch size of 1 and a maximum output length of 8192 tokens, using a challenging AIME25 problem where all three methods reached the token limit. Memory usage was profiled using torch.cuda.memory._record_memory_history.
Latency Breakdown. We report the decoding latency, throughput, and a detailed time breakdown for Standard, AlwaysThink, and TaH in Table 7. Here, Iter-1 forward and Iter-2 forward denote the total forward-pass time spent on the first and second latent iterations, respectively; Iter decider is the time for the iteration decider network to judge whether to continue iterating or verbalize; LoRA switching is the overhead of switching LoRA adapters; and Other includes tensor initialization, concatenation, and related bookkeeping.
Table 7: Per-component latency breakdown on a single A800 GPU.
Component Standard TaH AlwaysThink
Latency (s) Ratio(%) Latency (s) Ratio(%) Latency (s) Ratio(%)
Iter-1 Forward 210.6 100.0 229.8 76.2 224.1 30.0
Iter-2 Forward 29.6 9.8 384.7 51.5
Iter. Decider 10.5 3.5
LoRA Switching 7.5 2.5 91.1 12.2
Other 24.1 8.0 47.4 6.3
[…]
Table 8: Real-world decoding performance on a single A800 GPU, tested on AIME25 with 8K max token length. TPS denotes tokens per second.
Standard AlwaysThink TaH
Avg. Depth 1.00 2.00 1.06
Memory (GB) 4.3 6.8 4.6
Latency (s) 210.6 747.2 301.4
TPS 38.9 11.0 27.2
Discussion. We note that our current implementation is not yet optimized at the system level, so there remains room for further efficiency improvements. For example, the LoRA Switching and Other overheads (bookkeeping) are relatively high due to the Python-level implementation of dynamic control flow. These engineering optimizations are largely orthogonal to the algorithmic design of TaH, and we plan to continue refining the implementation to further reduce latency and memory overhead. The theoretical FLOPs and memory access analysis of TaH are provided in Appendix A.2.3.
A.2.3 Theoretical Efficiency Analysis
Following prior work Hoffmann et al. (2022); Yang et al. (2024); Ma et al. (2025), we analyze the computational and memory access overhead of TaH relative to the Standard and AlwaysThink baselines. Table 9 presents the average number of input/output tokens and latent iterations per token across five benchmarks. We use these statistics to calculate the theoretical computation and memory access costs for each method.
[…]
Table 9: Input tokens (shared across methods) and output token / iteration statistics for Standard, AlwaysThink, TaH, and TaH+ (Math-trained version).
Standard AlwaysThink TaH TaH+
Param. Dataset In. Out. Iter. Out. Iter. Out. Iter. Out. Iter.
0.6B AIME25 159 7450 1.00 7316 2.00 7648 1.05 7486 1.06
OlympiadBench 100 6599 1.00 6622 2.00 6631 1.09 6513 1.06
AMC23 85 6377 1.00 6368 2.00 6242 1.05 6145 1.05
MATH500 71 4823 1.00 5350 2.00 4877 1.05 4793 1.06
GSM8K 61 1955 1.00 2844 2.00 1923 1.07 1791 1.07
[…]
1.7B AIME25 159 7195 1.00 7173 2.00 7496 1.06 7498 1.06
OlympiadBench 100 6008 1.00 6484 2.00 6387 1.06 6258 1.06
AMC23 85 5681 1.00 7543 2.00 6122 1.04 5852 1.06
MATH500 71 4004 1.00 4414 2.00 4233 1.06 4286 1.06
GSM8K 61 1451 1.00 1644 2.00 1721 1.08 1686 1.08
[…]
Table 10: Decoding computation (GFLOPs) and memory access (GB) per output token for Standard, AlwaysThink, TaH and TaH+ methods.
Standard AlwaysThink TaH TaH+
Param. Dataset Comp. Mem. Comp. Mem. Comp. Mem. Comp. Mem.
0.6B AIME25 1.47 1.38 3.35 3.14 1.52 1.43 1.57 1.47
OlympiadBench 1.41 1.32 3.21 3.02 1.51 1.42 1.50 1.41
AMC23 1.40 1.31 3.17 2.97 1.43 1.34 1.46 1.37
MATH500 1.31 1.22 2.98 2.80 1.35 1.26 1.39 1.31
GSM8K 1.14 1.06 2.54 2.38 1.19 1.12 1.22 1.14
Average ratio $\times 1.00$$\times$ $\times 2.27$$\times$ $\times 1.04$$\times$ $\times 1.06$$\times$
1.7B AIME25 4.31 4.03 9.45 8.83 4.51 4.21 4.64 4.34
OlympiadBench 4.16 3.88 9.18 8.58 4.36 4.07 4.48 4.18
AMC23 4.12 3.85 9.54 8.91 4.24 3.96 4.43 4.13
MATH500 3.92 3.66 8.45 7.89 4.10 3.83 4.23 3.95
GSM8K 3.62 3.38 7.48 6.98 3.87 3.61 3.98 3.72
Average ratio $\times 1.00$$\times$ $\times 2.19$$\times$ $\times 1.05$$\times$ $\times 1.08$$\times$
A.2.4 Training Dynamics and Continuation Threshold
Figure 4(a) shows the stage 1 training dynamics, and Figure 4(b) reports GSM8K accuracy under different continuation thresholds (with iterate-twice ratios).
(a) Training dynamics of the LLM backbone on Qwen3-0.6B-Base. TaH converges rapidly and achieves lower perplexity.
(b) GSM8K accuracy with respect to continuation threshold. Numbers in brackets show the percentage of iterate-twice tokens.
Figure 4: Training dynamics and continuation threshold analysis.
A.2.5 Iteration Depth Beyond Two
@@
Hard Token Labeling. Previous works have proposed many methods to evaluate the hardness of each tokens in the training data, like through excess loss (Lin et al., 2024; Xie et al., 2023), entropy (Wang et al., 2025b; Chen et al., 2023) and prediction difference (Fu et al., 2025).
@@
Table 11: Performance comparison between TaH-2 and TaH-3 (maximum per-token iterations of 2 and 3, respectively). Iter.2 and Iter.3 denote the per-token percentages executing two and three iterations, respectively.
@@
A.2.6 Additional Latent Thinking Methods
Some latent thinking methods require pre-training and use base models other than Qwen3. We also compare with these methods, including Ponder (Zeng et al., 2025). Specifically, we adopt the released pretrained PonderingPythia-1.4B as the base model and perform SFT on the same training data. We observe that the fine-tuned model learns the stylistic patterns of the training data, but still underperforms substantially, which may be attributable to the limited capability of the PonderingPythia-1.4B backbone.
@@ A.2.6 Additional Latent Thinking Methods
A.2.7 Design Choice Exploration
We conduct controlled ablations where each variant changes a single component from the default TaH configuration to isolate its effect. Table 13 reports the final downstream results, and Figure 12 expands it by showing validation perplexity dynamics across different supervision signals and iteration policies. The naming convention matches Table 13. TaH with token-only supervision and the oracle policy yields lower perplexity than iter. decider and token+latent. Although the dynamic policy achieves the lowest perplexity, it fails on downstream tasks and often produces infinite-loop generations.
Table 13: Ablation study on design choices of TaH-0.6B. Each row varies one aspect from TaH configuration (marked with gray).
Variant MATH500 AMC23 Olympiad Average
TaH 51.2 32.5 23.9 35.9
Model Architecture
Iteration Depth (TaH: Neural decider)
Always-1 47.2 23.4 18.8 29.8//-6.1
Always-2 32.8 15.6 10.2 19.5//-16.4
Attention (TaH: Duo-causal)
Causal@iter1 47.8 24.4 19.4 30.5//-5.4
Causal@current 42.0 23.8 16.4 27.4//-8.5
Depth Adapters (TaH: LoRA + Residual)
w/o LoRA 51.6 29.7 22.4 34.6//-1.3
w/o LoRA & Res. 49.2 22.5 21.2 31.0//-4.9
Training Scheme
Superv. (TaH: Token-only)
Token+latent 49.4 29.6 15.9 31.6//-4.3
Iter. Policy (TaH: Oracle policy)
Decider-based 44.8 24.1 17.3 28.7//-7.2
Dynamic 11.0 2.8 2.7 5.5//-30.4
Discrepancy metric in $\pi$ (TaH: Top1 mismatch)
Cross-entropy 47.4 21.2 20.4 29.7//-6.2
Entropy 42.0 21.9 16.9 26.9//-9.0
Table 14: Impact of iteration schemes on Qwen3-0.6B (first 100 MATH500 samples).
Training Inference Accuracy
Standard Standard 52
AlwaysThink AlwaysThink 38//-14
AlwaysThink TaH-Oracle 77/+/+25
TaH-Oracle TaH-Decider 54/+/+02
TaH-Oracle TaH-Oracle 80/+/+28
A.2.8 Latent Overthinking
To analyze latent thinking patterns, we verbalize tokens from all iteration depths using their last-layer hidden states. The oracle method uses the oracle policy $\pi$ from Section 5.3 for iteration decision. (1) Generation. Since ground-truth tokens are unavailable during generation, we use predictions from the stronger DeepSeek-R1-Distill-Qwen-32B model (Guo et al., 2025) as proxy labels. Table 14 shows that the oracle policy substantially improves performance by verbalizing correct predictions immediately while iterating only on incorrect ones. With our trained iteration decider approximating the oracle, TaH outperforms both Standard and AlwaysThink baselines. However, the ideal oracle policy achieves even higher gains, indicating future potential. (2) Next-token prediction. We evaluate next-token prediction accuracy on the Open-R1 dataset, using the test model itself as the reference model in $\pi$. Figure 1 reveals that AlwaysThink produces more incorrect than correct revisions, demonstrating latent overthinking. In contrast, oracle-controlled iterations substantially increase correct revisions by selectively targeting hard tokens.
A.3 Additional Analysis
A.3.1 Oracle Policy and Hard Token Analysis
Metrics for Hard Token Labeling. We investigate different metrics for defining hard tokens to validate our choice of top-1 prediction mismatch. We compare three labeling strategies:
  1. 1.
    Top-1 Mismatch (TaH Default): Labels a token as hard if the reference model’s greedy prediction differs from the ground truth.
  2. 2.
    Entropy (TaH-Entropy): Labels a token as hard if the reference model’s prediction entropy exceeds a threshold.
  3. 3.
    Cross-Entropy (TaH-CE): Labels a token as hard if the reference model’s cross-entropy loss exceeds a threshold.
To ensure a fair comparison, for TaH-Entropy and TaH-CE, we set the thresholds such that the number of hard tokens in each sample matches the total ratio from the default Top-1 Mismatch policy. This isolates the impact of which tokens are selected, rather than how many.
Figure 5 compares the validation loss, and Table 15 reports downstream accuracy on 0.6B models. While cross-entropy (TaH-CE) improves over entropy labeling (TaH-Entropy), the Top-1 Mismatch policy (TaH) achieves superior performance across all benchmarks. This empirically verifies that directly targeting tokens where the model’s top-1 prediction is wrong is the most effective way to identify hard tokens for TaH.
Figure 5: Validation loss curves of 0.6B models trained with different oracle labeling metrics. The default Top-1 Mismatch policy yields the lowest validation loss.
Table 15: Performance comparison of different difficulty metrics (Entropy, Cross-Entropy, and Top-1 Accuracy) on 0.6B models. All methods mark the same total number of tokens as ”hard.”
Method MATH500 AMC23 OlympiadBench Average
TaH-Entropy 42.0 21.9 16.9 26.9
TaH-CE 47.4 21.2 20.4 29.7
TaH 51.2 32.5 23.9 35.9
Labeling Robustness. We investigate the robustness of hard-token labels with respect to the choice of reference model. We do so by analyzing the consistency of hard-token identification across different model scales (e.g., Qwen3-0.6B, 1.7B, and 4B).
First, we quantify the agreement between models. As shown in Figure 6, hard tokens exhibit high consistency across scales. Notably, even a smaller, less accurate reference model (1.7B) successfully identifies 81% of the hard tokens for a larger model (4B).
Second, to understand the quality of this overlap, we partition tokens into an overlap set (marked as hard by both models) and a non-overlap set (marked as hard by only one model). We plot the cross-entropy loss under each reference model in Figure 7. We observe that overlap tokens have substantially higher average cross-entropy ($\approx 2.0\times$) than non-overlap tokens for all reference models. This indicates that either reference model can identify this core set of ”hard” tokens, which corresponds to positions of genuine, high uncertainty. It reveals a consensus on hardness among models even of different sizes.
(a) Overlap between Qwen3-0.6B and 1.7B
(b) Overlap between Qwen3-1.7B and 4B
Figure 6: Venn diagrams illustrating the overlap of hard-token labels between different reference models. The high overlap proportions indicate that ”hard” tokens are largely consistent across model scales.
(a) Qwen3-1.7B: cross-entropy of overlap vs. non-overlap hard tokens (w.r.t. Qwen3-0.6B).
(b) Qwen3-0.6B: cross-entropy of overlap vs. non-overlap hard tokens (w.r.t. Qwen3-1.7B).
(c) Qwen3-4B: cross-entropy of overlap vs. non-overlap hard tokens (w.r.t. Qwen3-1.7B).
(d) Qwen3-1.7B: cross-entropy of overlap vs. non-overlap hard tokens (w.r.t. Qwen3-4B).
Figure 7: Token-level cross-entropy distributions of overlap and non-overlap hard tokens across different reference model pairs. For each pair of reference models (e.g., Qwen3-1.7B and Qwen3-0.6B), we plot the cross-entropy of tokens labeled as hard by both models (overlap) and by only one model (non-overlap) on both reference models.
A.3.2 Iteration Decider Robustness
We evaluate the iteration decider, trained on the general OpenR1 corpus, across three validation subsets (Math, Code, and QA) to quantify its robustness and cross-domain generalizability. As summarized in Table 16, the decider maintains high decision accuracy across all domains without any retraining.
Table 16: Iteration decider behavior and downstream gains on different validation subsets. The decider is trained once on general OpenR1 and evaluated without retraining.
Metric Math Code QA
Iteration Percentage 7.8% 10.7% 26.6%
Iteration Accuracy 86.7% 82.3% 76.6%
Benchmark Gain over Standard +6.8% +7.9% +5.8%
Despite being invoked on only 7.8-26.6% of tokens, the decider consistently yields 5.8-7.9% absolute accuracy gains over the standard single-pass baseline on all three domains. Moreover, the decider automatically adjusts its iteration rate according to task difficulty: it iterates more frequently on QA (26.6%) than on Math (7.8%), even under a fixed threshold $c_{\text{threshold}}=0.9$. This behavior indicates that the decider responds to intrinsic uncertainty signals in the model’s predictive distribution rather than memorizing domain-specific patterns, consistent with the token-level difficulty analysis in Appendix A.3.3.
A.3.3 Hard Token Identifiability
(a) Math
(b) Science
(c) Code
Figure 8: Output Logit entropy distribution at the first iteration of TaH, categorized by oracle policy’s difficulty labels (hard token) on the OpenR1 validation set (Math, QA, Code). The distinct separation between distributions confirms that TaH’s internal logits provide a strong, task-agnostic signal for identifying hard tokens.
Why is the iteration decider robust and generalizable across tasks? We investigate this by analyzing the intrinsic properties of ”hard” tokens.
We compute the token entropy of hard and easy tokens across three diverse subsets of the OpenR1 dataset (Math, Science, and Code). As shown in Figure 8, hard tokens exhibit a universal signature of significantly higher entropy ($5\times$) compared to easy tokens. This distinct separation confirms that ”hardness” is an intrinsic, robustly identifiable property of the model’s predictive state, rather than a complex, task-specific pattern. Given this clear signal, the neural iteration decider can easily learn reliable classification strategies that generalize well across different domains.
A.3.4 Token Alternation Pattern
$\Pr\big(c_{i}^{(1)}>c_{\text{threshold}}\mid t_{i}=t\big),$
But 34.3% So 13.63%
Wait 12.17%
Therefore 8.95%
So 17.7% So 28.17%
Therefore 13.67%
But 4.89%
A.3.5 Iteration Decision Error